OpenStudy (anonymous):

quadrilateral ABCD is a paralleogram in which angle B=100 degress. angle A+ angle C=___ degress.

OpenStudy (anonymous):

Opposite angles are congruent -> A=C, B=D, A+B+C+D=360.

OpenStudy (anonymous):

yeah i know that but i dont know what angle A and angle C is though

OpenStudy (anonymous):

You solve for them using the equation I gave you.

OpenStudy (anonymous):

but there is only one measurement so i need alittle more help then that

jhonyy9 (jhonyy9):

so how do you think if A=C and B=D but you know that B=D=100 degree so B+D=200 - and you know that A+B+C+D=360 so than A+C=360-200=160 so than A=C=160/2 =80 ok ?

OpenStudy (anonymous):

Do you know well the properties of parallelograms? And more importantly, do you know any basic algebra?

OpenStudy (anonymous):

okay